Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
taranto - 10:21 Monday 20 November 2023 (62500) Print this report
Comment to Restoring of the TCS configuration (62498)

After being called by the control room, at 8.30 UTC, the CO2 powers were checked. The NI DAS IN and OUT showed no issues, the WI CH was changed by 2%, while the NI CH by 14% in order to restore the CO2 actuators powers needed for CITF and lock acquisition. Moreover, at 09.03 UTC PR CHRoCC VDAC 0.08 --> 0.12 (3600 sec ramp time) by the operator.

Search Help
×

Warning

×